OpenStudy (mathmath333):

6 positive number are taken at random and multiplied together. Then what is the probability that the product ends in an odd digit other than 5.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ 6 positive number are taken at random and multiplied together.}\hspace{.33em}\\~\\ & \normalsize \text{ Then what is the probability that the product ends in an odd }\hspace{.33em}\\~\\ & \normalsize \text{ digit other than 5. }\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmate):

I assume "random" means that the probability of the last digit is uniformly distributed over the range 0-9. The resulting last digit depends only on the last digit of the six numbers, so without loss of generality (WLOG), we will consider only the product of six "random" digits. The resulting digit will be odd if and only if all six digits are odd. So what is the probability of getting 6 odd digits out of a uniform distribution? Out of the set {0,1,2,3,4,5,6,7,8,9}, there are 5 odd digits. So the probability of randomly selecting an odd digit is 5/10=1/2, resulting in an odd number. However, we know that 5 multiplied by any odd number will have 5 as a terminating digit, so we need to exclude 5 in ALL of the six numbers. That leaves us with 4, A={1,3,7,9} to choose from (out of 10). The probability of choosing any element of A from 10 digits is 4/10=2/5. The experiment is a 6 step experiment, each one independent of each other, so the multiplication rule applies. Can you figure out the probability of the final result (terminating with an odd digit excluding 5)?

OpenStudy (mathmath333):

2/5 ?

OpenStudy (mathmate):

It's 2/5 if you only pick one single number. Use the multiplication rule for 6 numbers that are multiplied together.

OpenStudy (mathmath333):

I still cant get ur hint

OpenStudy (dan815):

=]

OpenStudy (mathmath333):

i m depressed

OpenStudy (dan815):

hahahaha

OpenStudy (dan815):

im actually laughing irl rn

OpenStudy (dan815):

dang mathmate wrong something huge, can i just do it over

OpenStudy (dan815):

wrote*

OpenStudy (dan815):

6 numbers multiplied together no even, no 5 on the end

OpenStudy (mathmath333):

is it =2/5 *6

OpenStudy (dan815):

so no even numbers picked, and no 5 picked as that will gurantee a 5 or even number ending

OpenStudy (dan815):

the multiplication of all odd numbers is always off so u are safe there

OpenStudy (dan815):

odd*

OpenStudy (mathmath333):

answer was (0.4)^6

OpenStudy (mathmath333):

actuallyi m poor in english , cant totaly understand matemate paragraph

OpenStudy (dan815):

hence we have 4 choices for each of the places so 4^6/6!, as we dont care about the order u pick the 6 numbers so (4^6/6!) / (10^6/6!) = 0.4^6

OpenStudy (dan815):

the explaination is still same as above

OpenStudy (dan815):

what language do you speak?

OpenStudy (mathmath333):

rural marathi/hindi

OpenStudy (mathmath333):

हॅलो दान

OpenStudy (mathmath333):

https://www.youtube.com/watch?v=GHEX2RrY-cY

OpenStudy (dan815):

okay okay

OpenStudy (dan815):

do u want to start over?

OpenStudy (mathmath333):

how come \( (10^6/6!)\)

OpenStudy (dan815):

thats all the possible 6 digit products there are

OpenStudy (dan815):

this is another way to think about it uhh

OpenStudy (dan815):

u know we can pick only 4 digits out of the 10 digits

OpenStudy (dan815):

so (4/10)^6

OpenStudy (mathmath333):

but why ^{6}

OpenStudy (dan815):

for 6 digits

OpenStudy (dan815):